OpenStudy (anonymous):

(sin Θ − cos Θ)^2 + (sin Θ + cos Θ)^2

OpenStudy (anonymous):

Okay so I know that it will will be (sin Θ − cos Θ)^2 + 1 but i dont know what to do after

OpenStudy (anonymous):

the four answers are −sin2 Θ −cos2 Θ 0 2

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \rm sin Θ + cos Θ=1 }\) and \(\normalsize\color{blue}{ \rm (sin Θ - cos Θ)^2=sin^2Θ-2sinΘcosΘ+cosΘ= }\) \(\normalsize\color{blue}{ \rm 1-2sinΘcosΘ=1-sin(2Θ) }\) thus you get, \(\normalsize\color{blue}{ \rm 1-sin(2Θ)+1 }\) \(\normalsize\color{blue}{ \rm 2-sin(2Θ) }\)

OpenStudy (solomonzelman):

for the 2nd blue line it should be \(\normalsize\color{blue}{ \rm sin^2Θ-2sinΘcosΘ+cos^2(Θ) }\) at the end of the line.

OpenStudy (anonymous):

so what is the final answer?

OpenStudy (anonymous):

2-sin(2 Θ) is not in the answers

OpenStudy (vishweshshrimali5):

Here is the solution : \[\large{(\sin{\theta} - \cos{\theta})^2 + (\sin{\theta} + \cos{\theta})^2}\] \[\large{= \sin^2{\theta} + \cos^2{\theta} - 2\sin{\theta}\cos{\theta} + \sin^2{\theta} + \cos^2{\theta} + 2\sin{\theta}\cos{\theta}}\] Now, solve it after this. Remember that \(\large{\sin^2{\theta} + \cos^2{\theta} = 1}\)

OpenStudy (anonymous):

I don't know how to solve it

OpenStudy (anonymous):

can you please help me

OpenStudy (anonymous):

would the answer be 1?

OpenStudy (anonymous):

i mean 2

OpenStudy (vishweshshrimali5):

Yeah you are perfectly right.

OpenStudy (anonymous):

thank youuuu

OpenStudy (vishweshshrimali5):

:) np

OpenStudy (anonymous):

\[ (\sin (\theta )-\cos (\theta ))^2+(\sin (\theta )+\cos (\theta ))^2=\\\sin ^2(\theta )+\sin ^2(\theta )+\cos ^2(\theta )+\cos ^2(\theta )+2 \sin (\theta ) \cos (\theta )-2 \sin (\theta ) \cos (\theta )=\\2 \sin ^2(\theta )+2 \cos ^2(\theta )=2 \]